OpenStudy (anonymous):

How to solve general solution of the equation :( d square x / dt square) + 2dx/dt + 2x = 0. Help me please..

myininaya (myininaya):

x''+2x'+2=0

myininaya (myininaya):

solve the characteristic equation \[\lambda^2+2 \lambda+2=0 \text{ first } \]

myininaya (myininaya):

your solutions would be complex

myininaya (myininaya):

\[\lambda=a+bi \\ \text{ then solution is } \\ x=c_1 e^{at}\cos(bt)+c_2 e^{at} \sin(bt)\]

myininaya (myininaya):

type-o \[\lambda=a \pm bi\]

myininaya (myininaya):

@lyna9021 are you cool?

OpenStudy (anonymous):

Then the question ask to solve the equation given that t=0, x=1, and dx/dt=2

myininaya (myininaya):

ok so it wants us to also find those constants c1 and c2

OpenStudy (anonymous):

From the char equation, x I got 0.22 and -2.22

OpenStudy (anonymous):

Then what I need to do the next step? What is c1 and c2?

myininaya (myininaya):

how do you have that and leave in exact form

OpenStudy (anonymous):

take the derivative of x in the first post, you will have 2 equations with c1,c2, solve

myininaya (myininaya):

did you use the quadratic formula?

OpenStudy (anonymous):

fourth post*

OpenStudy (anonymous):

Yes

myininaya (myininaya):

you are going to leave in exact form and simplify as much as you can

myininaya (myininaya):

oops that is lambda not x

OpenStudy (anonymous):

Yes.I got answer is formula

myininaya (myininaya):

\[\lambda=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

Yes, then I got the answer -0.22 and 2.22

OpenStudy (anonymous):

Ohh that is for c1 n c2??

OpenStudy (anonymous):

c1,c2 are not the same as lambda

OpenStudy (anonymous):

you solve the characteristic equation in order to find c1,c2

OpenStudy (anonymous):

Okay.. What did the answer for c1 n c2?

OpenStudy (anonymous):

Okay.. What did the answer for c1 n c2?